OpenStudy (anonymous):

A uniform hollow cylinder is filled with equal mass amounts of pure water and oil such that the substances are completely separated if the pressure at the bottom of the water is 10.5 higher than it is at the top of the oil, then how high is the oil level above the bottom of the containter

OpenStudy (anonymous):

@phi @amistre64

OpenStudy (anonymous):

\[p_bot=p_0+p_{oil}+p_{h2o}\]

OpenStudy (anonymous):

@phi hmmm ican't figure this out... i know that the bottom will be the pressures of all these...

OpenStudy (anonymous):

and i know the masses are the same

OpenStudy (anonymous):

of oil and water

OpenStudy (phi):

I assume the top of the cylinder is open, and P_top is the pressure of the atmosphere If we make the cylinder have a base area of 1 cm^2 we can fill in the numbers. per http://www.engineeringtoolbox.com/liquids-densities-d_743.html olive oil has a density of about 0.9 g/ cm^3 and water has a density of 1 g/ cm^3 Can you make progress with this?

OpenStudy (anonymous):

i have 1000 kg/m^2 and 890kg/m^2

OpenStudy (anonymous):

should i first solve m=m

OpenStudy (phi):

solve m=m? I would find the force of the water: mass* g and then the pressure mass*g/base area of the cylinder ditto for the oil and you need the pressure of the atmosphere at sea level

OpenStudy (phi):

per wikipedia , a column of air one square centimeter in cross-section, measured from sea level to the top of the atmosphere, has a mass of about 1.03 kg

OpenStudy (anonymous):

i have no mas for water... i only know that mass of both liquids are thesame

OpenStudy (anonymous):

for force i need mass and i have no volume either

OpenStudy (anonymous):

so the force of the water is unable to be gotten, \[F_w=m_w*g\]

OpenStudy (anonymous):

the mass of water however is \[\rho V=\rho Ah\\]

OpenStudy (phi):

mass is density times volume density of water is 1000 kg/m^3 volume will be h1*1000 (assume base area is 1 m^2) mass of the oil is 890*h2 h1+h2= h

OpenStudy (anonymous):

\[\rho V=\rho Ah\]

OpenStudy (anonymous):

that's correct, we don't know the height of the cylinder so i'm still not seeing the point of this

OpenStudy (phi):

equal mass h1*1000 = h2*890

OpenStudy (anonymous):

\[i did this

OpenStudy (anonymous):

so i got something \[\frac{\rho_{h_2o}h_{h_2o}}{\rho_{oil}}\]

OpenStudy (anonymous):

i left rho's by themself because i ned to symbolically solve for my class

OpenStudy (anonymous):

actually we want the depth of the water correct

OpenStudy (phi):

mass of the water is 1000*h1= 1000*0.89*h2= 890 h2 mass of the oil is 890*h2 together 2*890*h2 so we have 1780 h2 + Ptop= 10.5*Ptop we need a number for Ptop

OpenStudy (anonymous):

\[\frac{\rho_{oil}h_{oil}}{\rho h_{h_2o}}=h_{h_2o}\]

OpenStudy (phi):

once we get any height we can solve for whatever they want...

OpenStudy (anonymous):

yes but wy not solve for the heigt of water because that will be the height frm the bottom or the container to the oil level

OpenStudy (phi):

yes but wy not solve for the heigt of water because that will be the height frm the bottom or the container to the oil level mostly because 0.89 is nicer than 1/0.89

OpenStudy (anonymous):

what idon't understand is when you did